# Curvature Tuning: Provable Training-free Model Steering From a Single Parameter

### v4

@@ 3.3 Curvature Tuning Operates as a Projection
This section provides a characterization of CT, by casting it as a projection of a ReLU network to a space of smooth functions. All proofs are deferred to Section C.1.
@@
For a ReLU network $f:\mathbb{R}^{d}\to\mathbb{R}$ with parameter $\mathbf{W}$ (collecting all weights and biases), for fixed $c\in{[}0,1{]}$ and $\beta\in{[}0,1)$, replacing every instance of ReLU with a CTU (Eq. 8 with hyperparameters $\beta,c$ is equivalent to projecting $f$ to a smooth function $f_{\beta,c}$ with bounded gradients and curvature, while keeping $\mathbf{W}$ fixed. Importantly, for $10<\beta<$ 1, $f_{\beta,c}$ enjoys higher local expressivity than $f$ for the same parameter $\mathbf{W}$, due to non-vanishing local curvature.
To conclude, we observe how varying $\beta$ modulates the curvature of the whole model function $f$ and, in turn, of the model’s decision boundaries. We begin by noting that for a deep network $f:\mathbb{R}^{d}\to\mathbb{R}^{k}$, the decision boundary between any class $i$ and $j$ is given by $\{\mathbf{x}\in\mathbb{R}^{d}:g(\mathbf{x}):=f_{i}(\mathbf{x})-f_{j}(\mathbf{x})=0\}$, for any $i,j=1,\ldots,k$ with $i\neq j$. Particularly, $g$ is itself a deep network, sharing the same parameters as $f$ up until the penultimate layer, after which the parameter is the vector $W_{i}^{L}-W_{j}^{L}$ and bias $\mathbf{b}^{L}_{i}-\mathbf{b}^{L}_{j}$. Importantly, when varying $\beta$ while keeping all model parameters fixed, the Jacobian $\nabla_{\mathbf{x}}g(\mathbf{x})$ and the Hessian $\nabla^{2}_{\mathbf{x}}g(\mathbf{x})$ are respectively given by the gradients and Hessian of $\mathbf{z}^{L-1}(\mathbf{x})$ – corresponding to the post-activation output of the $L-1$-th layer – weighted by $W^{L}_{i}-W^{L}_{j}$. Hence, modulating the non-linearity of activation functions via $\beta$ directly controls the curvature of both model function and its decision boundaries.  [^3] In the following, unless specified, we will thus refer interchangeably to the curvature of a DN mapping and that of its decision boundaries, whenever modulating non-linearities via CT.
Particularly, for $c=1$ (Eq. 8), as $\beta\to 0$, the activation becomes linear. Since modern DNs (e.g. MLP, CNN, RNN) are composed of activation functions interleaved with affine layers, it follows directly that the entire input-output mapping becomes affine when $\beta\to 0$. In this setting, the curvature of the mapping—defined as the norm of its Hessian—becomes zero. As a result, transitioning from the original DN mapping ($\beta=1$) to the linear setting effectively modulates the network decision boundary curvature, reducing it continuously to zero in the limit. For $c<1$, the model retains non-vanishing local curvature, while the mapping becomes smooth.
@@ 4 Enhancing Model Generalization and Robustness with CT
In this section, we empirically validate the effectiveness of CT across multiple settings. We first show that CT improves generalization (Section 4.1) and enhances robustness (Section 4.2). Section 4.3 shows that Trainable CT achieves gains comparable to LoRA with fewer parameters. Finally, we demonstrate CT’s effectiveness on transformers despite only partial guarantees (Section 4.4). GPU and seed details are provided in Appendix B.
4.1 Improving generalization on downstream datasets with CT
In this subsection, we evaluate the effectiveness of CT in improving model generalization on a variety of downstream datasets. Specifically, we transfer ImageNet-pretrained ResNet-18, ResNet-50, and ResNet-152 models to 12 downstream tasks, including Arabic Characters [29], Arabic Digits [30], Beans [31], CUB-200-2011 [32], DTD [33], FashionMNIST [34], FGVC-Aircraft [35], Flowers102 [36], Food101 [37], and three subsets from MedMNIST-PathMNIST, OCTMNIST, and DermaMNIST [38]. Each dataset is split into training, validation, and test sets, with details in Section B.1.
To apply CT, we replace all ReLUs in the backbone with CTUs, freeze the model, and train a linear classifier on the penultimate layer. The optimal $\beta$ is selected via grid search over $\beta\in[0.7,1]$ (step size 0.01) using validation accuracy, and the corresponding test accuracy is reported. For the baseline, we train a linear classifier on the frozen original model and report the test accuracy of the checkpoint that performs best on the validation set. All linear classifiers use the same training setup detailed in Section B.1.
Table 1: Mean accuracy (%) over three runs of ImageNet-pretrained ResNet-18/50 when transferred to 12 downstream datasets. The second row under each method indicates the number of trainable parameters (excluding the linear classifier). CT outperforms linear probing on the frozen backbone, and Trainable CT surpasses LoRA (rank 1). Results for ResNet-152 can be found in Appendix Table 4.
@@ 4.1 Improving generalization on downstream datasets with CT
Dataset Frozen CT LoRA Train CT Frozen CT LoRA Train CT
@@
Table 1 and Appendix Table 4 show mean accuracy over three runs for linear probing with and without CT on 12 downstream datasets using ResNet-18/50/152 backbones. CT consistently improves generalization, with average relative gains of 1.97%, 1.16%, and 0.02% respectively.
We also report the average optimal $\beta$ across datasets: 0.84 for ResNet-18, 0.94 for ResNet-50, and 0.96 for ResNet-152 (full results in Appendix Table 5). These values are consistently close to 1, suggesting the search range can be narrowed for efficiency. The upward trend with model size indicates that larger models require less curvature adjustment, which is intuitive as deeper networks can approximate complex curvature more effectively. Example accuracy curves in Appendix Fig. 4 show that accuracy varies smoothly with $\beta$ and typically peaks in the middle of the search range.
4.2 Improving robustness on adversarial and corrupted data with CT
@@
In this subsection, we demonstrate that CT can enhance model robustness using benchmarks from RobustBench [14]. We evaluate robustness of ResNet-18, ResNet-50 and ResNet-152 on CIFAR-10, CIFAR-100, and ImageNet using the official $\ell_{2}$ and $\ell_{\infty}$ adversarial benchmarks from RobustBench [14], as well as the common corruption benchmark [39], which contains naturally perturbed images. For specific evaluation configurations, see Section B.2.
It is worth noting that here we apply CT in a straightforward manner to assess its robustness benefits, without focusing on how to find the optimal $\beta$ value. Specifically, we replace all ReLU activations in the backbone with CTUs and perform a grid search over $\beta\in[0.7,1]$ with a step size of 0.01. We report the value that achieves the best performance on each benchmark.
@@
Table 2: Mean robust accuracy (%) over three runs of ImageNet-pretrained ResNet-18/50/152 under $\ell_{2}$$\ell_{\infty}$ attacks and corruptions on CIFAR-10/100 and ImageNet. CT yields substantial improvements under $\ell_{\infty}$ attacks, moderate gains under $\ell_{2}$ attacks, and marginal improvements under common corruptions, with the selected $\beta$ values generally close to 1.
@@
Model Dataset Base CT $\beta$ Base CT $\beta$ Base CT $\beta$
@@
As summarized in Table 2, CT is particularly effective against $\ell_{\infty}$ attacks, achieving large relative improvements of 44.01%, 1032.64%, and 1494.46% for ResNet-18, ResNet-50, and ResNet-152, respectively. [^4] We exclude the ResNet-18 result on ImageNet under $\ell_{\infty}$ from this computation since the baseline robust accuracy is 0. The improvements under $\ell_{2}$ attacks are moderate—1.65%, 3.62%, and 0.39%—while the gains on corruption robustness are marginal. The corresponding average optimal $\beta$ values, also shown in Table 2, are consistently close to 1. This suggests that modest curvature modulation is sufficient to achieve improved robustness, showing again the efficiency of CT in practice.
4.3 Trainable CT is comparable to LoRA
In this subsection, we show that Trainable CT further improves generalization, achieving performance comparable to LoRA with fewer parameters. We conduct experiments using the same setup as in Section 4.1, evaluating both Trainable CT and LoRA. For both methods, we replace the original linear classifier in each model with an appropriate one, and apply the respective method to the pretrained backbone. In Trainable CT, we initialize all $\beta$ parameters to 0.8 and all $c$ parameters to 0.5. For LoRA, we set the rank $r=1$ and scale $\alpha=1$, focusing on the comparison rather than tuning LoRA’s hyperparameters. We apply LoRA to all convolutional and linear layers in the backbone, and provide implementation details in Appendix E. Refer to Section B.3 for training configurations we apply for the two methods. Here we report the test accuracy of the best checkpoint on the validation set.
The results, summarized in Table 1 and Appendix Table 4, show that Trainable CT achieves the highest performance across all methods, with average relative improvements on ResNet-18, 50, and 152 of 6.75%, 8.59%, and 8.34% over the baseline; 4.62%, 7.14%, and 8.51% over CT; and 10.20%, 4.64%, and 1.70% over LoRA. Importantly, Trainable CT achieves better performance than LoRA with fewer parameters. As reported in Table 1 and Appendix Table 4, the number of trainable parameters (excluding the classifier) used by Trainable CT amounts to only 11.05%, 57.20%, and 59.09% of that used by LoRA on ResNet-18, 50, and 152, respectively—even with LoRA operating at its lowest-rank setting ($r=1$). This highlights the parameter efficiency of our approach.
To better understand how Trainable CT behaves during training, we analyze the distributions of learned $\beta$ and $c$ values (full statistics provided in Appendix Tables 6 and 7). We observe a high degree of within-model variation, with standard deviations ranging from 0.31 to 0.38, while the means remain remarkably stable across architectures: 0.69 to 0.74 for $\beta$ and 0.57 to 0.59 for $c$. These mean values are close to those used in CT, though the learned $\beta$ values tend to be smaller than the optimal shared $\beta$ found in CT (0.84 to 0.96), while the learned $c$ values are larger than the fixed $c=0.5$.
We further visualize the distributions in Appendix Figs. 5 and 6. In most datasets, as shown in Appendix Fig. 5 (OCTMNIST), both $\beta$ and $c$ exhibit a sharp U-shaped distribution—concentrating near 0 and 1 with a flat middle. This suggests that Trainable CT leverages its parameter flexibility to assign values at the extremes, producing an effective average close to the manually chosen settings in CT, rather than concentrating around the mean values themselves. [^5] This behavior may in part be influenced by the sigmoid-based parameterization used in our implementation of Trainable CT to constrain $\beta$ and $c$ during training. In a few datasets, we observe deviations from this trend, as exemplified in Appendix Fig. 6 (DTD). Nonetheless, a consistent pattern is that for any given dataset, the distributions remain visually similar across all models.
4.4 CT shows promise on transformers and emerging architectures
@@
In this subsection, we investigate the effectiveness of CT on transformer architectures. Unlike ResNets, transformers include attention layers that fall outside the max-affine spline framework, weakening theoretical guarantees. However, the feed-forward blocks with convex, piecewise-affine activations (e.g., ReLU) still retain them.
While our method is naturally suited for ReLU activations, most transformers use activation functions such as GELU. To bridge this gap, we replace all GELUs in Swin-T and Swin-S with ReLU, and pretrain the models from scratch on Imagenette (details can be found in Section B.4). Then as before, we transfer the models to the same 12 downstream datasets and apply all four methods: the frozen baseline, CT, LoRA, and Trainable CT. The transfer protocols remain identical to those in Section 4.1 and Section 4.3, except that LoRA is now applied to all QKV projection layers.
@@
Table 3: Mean accuracy (%) over three runs of Imagenette-pretrained Swin-T/S when transferred to 12 downstream datasets. The second row under each method indicates the number of trainable parameters (excluding the linear classifier). CT outperforms linear probing on Swin-T, while Trainable CT significantly improves over linear probing on both models but underperforms LoRA.
@@
Dataset Frozen CT LoRA Train CT Frozen CT LoRA Train CT
(0) (1) (74832) (532) (0) (1) (148560) (868)
Arabic Characters 30.67 31.08 56.32 41.95 31.81 31.16 62.16 40.88
Arabic Digits 83.71 85.24 97.54 90.82 80.74 81.11 97.91 91.44
Beans 60.68 61.46 75.52 68.49 55.99 54.43 73.96 67.71
CUB-200 4.82 4.87 7.42 6.09 4.46 4.02 9.19 6.71
DTD 15.92 15.90 16.99 17.04 16.03 15.78 18.67 17.66
FashionMNIST 73.81 74.01 83.90 77.07 73.28 73.29 86.15 75.76
FGVC-Aircraft 4.57 4.47 5.59 6.14 4.61 4.74 6.55 6.16
Flowers102 14.09 14.01 16.66 16.53 12.93 13.12 17.28 17.99
Food101 14.85 14.79 18.17 15.20 14.22 14.28 19.41 14.50
DermaMNIST 70.24 70.99 74.08 71.37 69.23 70.34 73.93 70.59
OCTMNIST 49.60 51.37 63.53 53.23 48.07 47.93 63.90 51.23
PathMNIST 76.73 77.78 81.31 77.35 74.82 76.54 76.62 78.59
Average 41.64 42.16 49.75 45.11 40.52 40.56 50.48 44.94
The results, shown in Table 3, indicate that CT yields a modest average relative improvement of 0.68% over the frozen baseline on Swin-T, but slightly underperforms the baseline by 0.58% on Swin-S. In contrast, Trainable CT delivers substantial improvements, outperforming the baseline by 13.32% and 17.93% on Swin-T and Swin-S. Despite these gains, Trainable CT underperforms LoRA in this setting, trailing by 8.29% and 11.89% on Swin-T and Swin-S. We attribute this performance gap to two main factors. First, as shown in Table 3, Trainable CT introduces only 0.71% and 0.58% as many trainable parameters as LoRA on Swin-T and Swin-S—a much lower ratio than in the ResNet experiments. Second, the pretraining dataset, Imagenette, is relatively small, containing only 10 classes and 9,469 training samples. This limited scale may lead to underpowered representations for downstream transfer, as reflected in the overall low baseline accuracy across datasets. In such settings, LoRA’s higher capacity may allow it to more effectively adapt the pretrained model to diverse downstream tasks. Thus though Trainable CT is outperformed by LoRA in this scenario, the results still demonstrate the potential of CT in transformer models.
Moreover, recent work advocating a return to ReLU in transformers [40] suggests that CT can be applied to a broader class of state-of-the-art architectures with minimal modification. Finally, although our experiments focus on ReLU-based networks, CT is readily extensible to architectures using SiLU or Softplus activations, as both are special cases of the general CTU formulation in Eq. 8. This includes emerging models like Mamba [41], which uses SiLU.

### v5

@@ 3.3 Curvature Tuning operates as a projection
This section provides a characterization of CT, by casting it as a projection of a ReLU network to a space of smooth functions. All proofs are deferred to Section˜C.1.
@@
For a ReLU network $f:\mathbb{R}^{D}\to\mathbb{R}$ with parameter $\mathbf{W}$ (collecting all weights and biases), for fixed $c\in{[}0,1{]}$ and $\beta\in{[}0,1)$, replacing every instance of ReLU with a CTU (Eq.˜8 with hyperparameters $\beta,c$ is equivalent to projecting $f$ to a smooth function $f_{\beta,c}$ with bounded gradients and curvature, while keeping $\mathbf{W}$ fixed. Importantly, for $0<\beta<1$, $f_{\beta,c}$ enjoys higher local expressivity than $f$ for the same parameter $\mathbf{W}$, due to non-vanishing local curvature.
To conclude, we observe how varying $\beta$ modulates the curvature of the whole model function $f$ and, in turn, of the model’s decision boundaries. We begin by noting that for a deep network $f:\mathbb{R}^{D}\to\mathbb{R}^{K}$, the decision boundary between any class $i$ and $j$ is given by $\{\mathbf{x}\in\mathbb{R}^{D}:g(\mathbf{x}):=f_{i}(\mathbf{x})-f_{j}(\mathbf{x})=0\}$, for any $i,j=1,\ldots,K$ with $i\neq j$. Particularly, $g$ is itself a deep network, sharing the same parameters as $f$ up until the penultimate layer, after which the parameters are the vector $W_{i}^{L}-W_{j}^{L}$ and the bias $\mathbf{b}^{L}_{i}-\mathbf{b}^{L}_{j}$. Importantly, when varying $\beta$ while keeping all model parameters fixed, the Jacobian $\nabla_{\mathbf{x}}g(\mathbf{x})$ and the Hessian $\nabla^{2}_{\mathbf{x}}g(\mathbf{x})$ are respectively given by the gradients and Hessian of $\mathbf{z}^{L-1}(\mathbf{x})$ – corresponding to the post-activation output of the L1L-1-th layer—weighted by $W^{L}_{i}-W^{L}_{j}$. Hence, modulating the nonlinearity of activation functions via $\beta$ directly controls the curvature of both model function and its decision boundaries. [^4] In the following, unless specified, we will thus refer interchangeably to the curvature of a DN mapping and that of its decision boundaries whenever modulating nonlinearities via CT.
Particularly, for $c=1$ (Eq.˜8), as $\beta\to 0$, the activation becomes linear. Since modern DNs (e.g. MLP, CNN, RNN) are composed of activation functions interleaved with affine layers, it follows directly that the entire input-output mapping becomes affine when $\beta\to 0$. In this setting, the curvature of the mapping—defined as the norm of its Hessian—becomes zero. As a result, transitioning from the original DN mapping ($\beta=1$) to the linear setting effectively modulates the network decision boundary curvature, reducing it continuously to zero in the limit. For $c<1$, as $\beta\to 0$, the model retains non-vanishing local curvature, while the mapping becomes smooth.
@@ 4 Enhancing model generalization and robustness with CT
In this section, we empirically validate the effectiveness of CT across multiple settings. We first demonstrate that S-CT improves generalization (Section˜4.1), while T-CT achieves improvement comparable to LoRA with substantially fewer parameters (Section˜4.2). Then we show that CT can improve the robustness of models through its implicit bias (Section˜4.3). Finally, we demonstrate CT’s effectiveness on transformers despite only partial guarantees (Section˜4.4). GPU and seed details are provided in Appendix˜B.
4.1 Improving generalization on downstream datasets with S-CT
In this subsection, we evaluate the effectiveness of S-CT in improving model generalization on a variety of downstream datasets. Specifically, we transfer ImageNet-pretrained ResNet-18, ResNet-50, ResNet-152 and VGG-11 models to 12 downstream tasks, including Arabic Characters [12], Arabic Digits [11], Beans [28], CUB-200-2011 [37], DTD [7], FashionMNIST [38], FGVC-Aircraft [27], Flowers102 [31], Food101 [3], and three subsets from MedMNIST—PathMNIST, OCTMNIST, and DermaMNIST [39]. Each dataset is split into training/validation/test sets (details in Section˜B.1).
To apply S-CT, we replace all ReLUs in the backbone with CTUs, freeze the model, and train a linear classifier on the penultimate layer. The optimal $\beta$ is selected via grid search over $\beta\in[0.7,1]$ (step size 0.01) using validation accuracy, and the corresponding test accuracy is reported. For the baseline, we train a linear classifier on the frozen original model and report the test accuracy of the checkpoint that performs best on the validation set. All linear classifiers use the same training setup detailed in Section˜B.1.
Table 1: Mean accuracy (%) over three runs of ImageNet-pretrained ResNet-18/50 when transferred to 12 downstream datasets. The second row under each method indicates the number of trainable parameters (excluding the linear classifier). S-CT outperforms linear probing on the frozen backbone, and T-CT surpasses LoRA (rank 1). Full results (± std) in Table˜S1.
@@ 4.1 Improving generalization on downstream datasets with S-CT
Dataset Frozen S-CT LoRA T-CT Frozen S-CT LoRA T-CT
@@
Tables˜1 and S1 show mean accuracy over three runs for linear probing with and without S-CT on 12 downstream datasets using ResNet-18/50/152 and VGG-11 backbones. S-CT consistently improves generalization, with average relative gains of 1.97%, 1.16%, 0.02%, and 0.71% respectively. We also report the average optimal $\beta$ across datasets: 0.84 for ResNet-18, 0.94 for ResNet-50, 0.96 for ResNet-152 and 0.90 for VGG-11 (full results in Table˜S3). These values are consistently close to 1, suggesting the search range can be narrowed for efficiency. Example accuracy curves in Fig.˜S2 show that accuracy varies smoothly with $\beta$ and typically peaks in the middle of the search range.
Additionally, we conduct ablation experiments on the choice of $c=0.5$ for S-CT under the same settings. As shown in Table˜S4, setting $c=0.5$ yields better performance than $c=1$ (SiLU) or $c=0$ (Softplus) by 1.99% and 0.56% respectively on ResNet-18, and by 0.94% and 0.57% on ResNet-50, while performing slightly worse by 0.05% and 0.73% on ResNet-152.
4.2 T-CT is comparable to LoRA with fewer parameters
In this subsection, we show that T-CT further improves generalization, achieving performance comparable to LoRA with fewer parameters. We conduct experiments using the same setup as in Section˜4.1, evaluating both T-CT and LoRA. For both methods, we replace the original linear classifier in each model with an appropriate one, and apply the respective method to the pretrained backbone. In T-CT, we initialize all $\beta$ parameters to 0.8 and all $c$ parameters to 0.5. For LoRA, we apply it to all convolutional and linear layers in the backbone (implementation details in Appendix˜E). And we set the rank $r=1$ and scale $\alpha=1$, so that it has a comparable number of parameters to T-CT. Refer to Section˜B.2 for training configurations applied to the two methods. Here we report the test accuracy of the best checkpoint on the validation set.
The results, summarized in Tables˜1 and S1, show that T-CT achieves the best performance across all methods, with average relative improvements on ResNet-18/50/152 and VGG-11 of 6.75%, 8.59%, 8.34% and 5.53% over the baseline; 4.62%, 7.13%, 8.35% and 4.73% over S-CT; and 10.20%, 4.64%, 1.70% and 4.05% over LoRA. Importantly, T-CT achieves better performance than LoRA with fewer parameters. As reported in Tables˜1 and S1, the number of trainable parameters (excluding the classifier) used by T-CT amounts to only 11.05%, 57.20%, 59.09% and 36.39% of that used by LoRA on the four models respectively—even with LoRA operating at its lowest-rank setting ($r=1$). This highlights the parameter efficiency of our approach. We further compare T-CT and LoRA with varying ranks ($r\in\{1,2,4\}$) and scaling factors ($\alpha\in\{r,2r,4r\}$) in Table˜S5. For each dataset and LoRA rank, we report the best test accuracy achieved among the candidate scaling factors. As shown, T-CT can still outperform LoRA under this more challenging setting, achieving relative improvements of 9.65%, 6.56%, and 3.34% over the three ranks on ResNet-18; 4.18%, 2.27%, and 0.19% on ResNet-50; and 1.22%, –0.91%, and –1.84% on ResNet-152, further demonstrating the effectiveness and parameter efficiency of T-CT.
To better understand how T-CT behaves during training, we analyze the distributions of learned $\beta$ and $c$ values (full statistics provided in Tables˜S6 and S7). We observe a high degree of within-model variation, with standard deviations ranging from 0.31 to 0.38, while the means remain remarkably stable across architectures: 0.69 to 0.79 for $\beta$ and 0.57 to 0.61 for $c$. These mean values are close to those used in S-CT, though the learned $\beta$ values tend to be smaller than the optimal shared $\beta$ found in S-CT (0.84 to 0.96), while the learned $c$ values are larger than the fixed $c=0.5$. We further visualize the distributions in Figs.˜S3 and S4. In most datasets, as shown in Fig.˜S3 (OCTMNIST), both $\beta$ and $c$ exhibit a sharp U-shaped distribution—concentrating near 0 and 1 with a flat middle. This suggests that T-CT leverages its parameter flexibility to assign values at the extremes, producing an effective average close to the manually chosen settings in S-CT, rather than concentrating around the mean values themselves. [^5] This behavior may in part be influenced by the sigmoid-based parameterization used in our implementation of T-CT to constrain $\beta$ and $c$ during training. In a few datasets, we observe deviations from this trend, as exemplified in Fig.˜S4 (DTD). Nonetheless, a consistent pattern is that for any given dataset, the distributions remain visually similar across models.
4.3 Improving model robustness through CT’s implicit bias
@@
In this subsection, we demonstrate that CT exhibits an implicit bias toward enhancing model robustness, using benchmarks from RobustBench [8]. We evaluate the robustness of ResNet-18/50/152 on CIFAR-10/100 and ImageNet using the official $\ell_{2}$ and $\ell_{\infty}$ adversarial benchmarks from RobustBench [8] and the common corruption benchmark [17]. See Section˜B.3 for more details.
More specifically, we first show that S-CT can improve robustness without any robustness-oriented objective, by applying it to pretrained models, performing a grid search over $\beta\in[0.7,1]$ with a step size of 0.01, and reporting the best robust accuracy achieved. We then show that T-CT, when used for finetuning (from ImageNet to CIFAR-10/100 following the same training configurations as in Sections˜4.1 and 4.2), also enhances robustness despite the absence of a robustness objective—the finetuned models achieve stronger robustness than those trained with linear probing or LoRA ($r=1$).
@@
Table 2: Mean robust accuracy (%) over three runs of ImageNet-pretrained ResNet-18/50/152 under $\ell_{2}$$\ell_{\infty}$ attacks and corruptions on CIFAR-10/100 and ImageNet. S-CT yields substantial improvements under $\ell_{\infty}$ attacks, with the selected $\beta$ values close to 1. Full results (± std) in Table˜S8.
@@
Model Dataset Frozen S-CT $\beta$ Frozen S-CT $\beta$ Frozen S-CT $\beta$
@@
Table 3: Mean robust accuracy (%) over three runs of ImageNet-pretrained ResNet-18/50/152 transferred to CIFAR-10/100 under $\ell_{2}$, $\ell_{\infty}$ attacks, and corruptions. T-CT improves $\ell_{\infty}$ robustness significantly compared to linear probing and LoRA. Full results (± std) in Table˜S9.
$\ell_{2}$ $\ell_{\infty}$ Corruption
Model Dataset Frozen LoRA T-CT Frozen LoRA T-CT Frozen LoRA T-CT
ResNet18 CIFAR10 8.47 5.93 8.93 0.30 0.70 1.57 21.34 13.59 16.83
CIFAR100 1.57 0.77 1.10 0.03 0.07 0.17 5.10 2.96 4.62
Average 5.02 3.35 5.01 0.16 0.38 0.87 13.22 8.28 10.72
ResNet50 CIFAR10 6.23 4.57 6.83 0.20 0.33 2.43 16.23 11.69 12.68
CIFAR100 0.70 0.37 0.47 0.00 0.03 0.07 3.47 2.04 1.61
Average 3.47 2.47 3.65 0.10 0.18 1.25 9.85 6.86 7.14
ResNet152 CIFAR10 8.03 4.63 8.00 0.43 0.20 5.10 13.82 11.33 9.83
CIFAR100 0.90 0.47 0.50 0.17 0.00 0.00 2.07 2.13 1.72
Average 4.46 2.55 4.25 0.30 0.10 2.55 7.94 6.73 5.78
@@
For S-CT, as summarized in Table˜2, it is particularly effective against $\ell_{\infty}$ attacks, achieving substantial relative improvements of 44.01%, 1032.64%, and 1494.46% for ResNet-18, ResNet-50, and ResNet-152.666We exclude from relative-improvement computation any cases where the baseline robust accuracy is 0. Improvements under $\ell_{2}$ attacks and common corruptions are comparatively moderate. The corresponding optimal $\beta$ values, also shown in Table˜2, are consistently close to 1, suggesting that modest curvature modulation suffices to improve robustness—further highlighting the practical efficiency of S-CT. For T-CT, as shown in Table˜3, it likewise enhances $\ell_{\infty}$ robustness significantly, achieving average relative improvements over linear probing of 445.00%, 1115.00%, and 493.02%, and over LoRA of 133.57%, 384.85%, and 2450.00%. These results collectively demonstrate CT’s implicit bias toward robustness enhancement, further validating its effectiveness beyond standard generalization improvements.
4.4 CT shows promise on transformers
@@
In this subsection, we investigate the effectiveness of T-CT in improving the generalization of transformer architectures. Unlike ResNets, transformers include attention layers that fall outside the max-affine spline framework and typically employ non-ReLU activation functions (e.g., GELU), which weakens our theoretical guarantees.
Concretely, we apply T-CT to ImageNet-pretrained Swin-T and Swin-S models, both of which use GELU activations in their feed-forward blocks. As before, we transfer these models to the same 12 downstream datasets and compare T-CT against linear probing on the frozen backbone and LoRA ($r=1$, $\alpha=1$). Additional training details are provided in Section˜B.4.
@@
Table 4: Mean accuracy (%) over three runs of ImageNet-pretrained Swin-T/S when transferred to 12 downstream datasets. The second row under each method indicates the number of trainable parameters (excluding the linear classifier). T-CT improves over linear probing but underperforms LoRA. Full results (± std) in Table˜S10.
@@
Dataset Frozen LoRA T-CT Frozen LoRA T-CT
(0) (74832) (532) (0) (148560) (868)
Arabic Characters 83.48 93.24 85.02 83.83 94.38 86.65
Arabic Digits 98.14 99.19 98.47 98.28 99.19 98.39
Beans 88.28 94.01 89.06 90.89 95.05 91.41
CUB-200 73.42 78.73 74.33 72.66 79.45 73.40
DTD 70.66 70.99 71.45 69.77 71.56 72.43
FashionMNIST 89.89 93.15 90.23 89.75 93.52 89.85
FGVC-Aircraft 48.06 48.29 47.58 44.36 51.94 45.72
Flowers102 86.66 90.22 85.35 83.24 87.67 85.08
Food101 77.05 83.69 78.90 77.59 85.17 79.45
DermaMNIST 75.83 76.71 75.86 76.64 78.15 77.14
OCTMNIST 69.97 76.30 67.97 66.90 76.97 69.07
PathMNIST 89.14 92.26 91.73 89.74 92.79 92.13
Average 79.22 83.06 79.66 78.64 83.82 80.06
The results, presented in Table˜4, show that T-CT yields average relative improvements over linear probing of 0.48% and 1.94% on Swin-T and Swin-S, respectively. However, it underperforms LoRA in this setting, trailing by 4.01% and 4.76% on the two models. Nonetheless, as shown in Table˜4, T-CT requires only 0.71% and 0.58% as many trainable parameters as LoRA on Swin-T and Swin-S—a much lower ratio than in the ResNet experiments. Thus, despite its lower performance, the results still highlight its potential for transformer architectures. It is also worth noting that in the current implementation, CT is applied only to the feed-forward blocks, which constitute a relatively small portion of the transformer, while the attention layers make up the majority. Extending CT to modulate the curvature within attention mechanisms—such as by tuning the temperature in the softmax of the attention block—is an interesting future work.